OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (callisto):

let x and y be the no. of adults and students respectively x+y=150 => y=150-x -(1) 45x+28.5y = 4770 =>45x+28.5y = 4770 -(2) put (1) into (2) 45x +28.5 (150-x ) = 4770 16.5x = 495 x=30 therefore no. of students = 150 - 30 = 120 therefore no of adults = 30 and no of students = 120

OpenStudy (anonymous):

Thank You.. can you help me with one more please

OpenStudy (callisto):

the last one :)